OpenStudy (anonymous):

Please help me? (: Solve x + 3y = 93x - 3y = -13 (1, 10/3) (1, -10/3) (-1, 3/10) (1, -3/10)

jimthompson5910 (jim_thompson5910):

where does the first equation end and the second one begin?

OpenStudy (anonymous):

It's combined

jimthompson5910 (jim_thompson5910):

Is it x + 3y = 9 3x - 3y = -13 or is it x + 3y = 93 x - 3y = -13

OpenStudy (anonymous):

I believe the first one

OpenStudy (anonymous):

WAIT, it's the second

OpenStudy (anonymous):

x + 3y = 93x-3y

jimthompson5910 (jim_thompson5910):

x + 3y = 93 x - 3y = -13 ------------ 2x+0y = 80 So 2x = 80 what is x?

OpenStudy (anonymous):

( I know it's confusing, but that's how my professor has me do it)

OpenStudy (anonymous):

I don't know, i suppose I should find out. I think it has you wanting to solve for x.

jimthompson5910 (jim_thompson5910):

hmm it can't be the second one, none of the answer choices match

OpenStudy (anonymous):

It may have me solving for both x and y

jimthompson5910 (jim_thompson5910):

x + 3y = 9 3x - 3y = -13 -------------- 4x + 0y = -4 so 4x = -4 what is x?

OpenStudy (anonymous):

Try it the way I mentioned :)

jimthompson5910 (jim_thompson5910):

What do you get when you solve 4x = -4 for x?

OpenStudy (anonymous):

-16

OpenStudy (anonymous):

But that's irrelevant

jimthompson5910 (jim_thompson5910):

no, you don't multiply

jimthompson5910 (jim_thompson5910):

4 times something = -4 what is that "something"?

OpenStudy (anonymous):

I have no clue

jimthompson5910 (jim_thompson5910):

well 4 times 1 = 4 right?

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

so x = -1

OpenStudy (anonymous):

:) oh

jimthompson5910 (jim_thompson5910):

now use this to find y

OpenStudy (anonymous):

Ok, hmmm

OpenStudy (anonymous):

Point me in the right direction?

OpenStudy (anonymous):

I don't know if it'd be 10 over 3 or the opposite

jimthompson5910 (jim_thompson5910):

if x = -1, then x + 3y = 9 -1 + 3y = 9 3y = 9+1 3y = 10 What's the last step?

OpenStudy (anonymous):

to figure out why

OpenStudy (anonymous):

To figure out what "Y' equals

jimthompson5910 (jim_thompson5910):

exactly, how do we do that?

OpenStudy (anonymous):

10+3 = 13

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

get y by itself

jimthompson5910 (jim_thompson5910):

you need to undo the multiplication by doing what?

OpenStudy (anonymous):

or divide

OpenStudy (anonymous):

on both sides?

jimthompson5910 (jim_thompson5910):

yes, divide both sides by 3

OpenStudy (anonymous):

10/3 3.3333333

jimthompson5910 (jim_thompson5910):

i'd keep it as a fraction, so y = 10/3

OpenStudy (anonymous):

oh my god youre the best tutor

jimthompson5910 (jim_thompson5910):

thanks lol

OpenStudy (anonymous):

But one quick question, how can you decipher whether it'd be 10/3 or 3/10?

jimthompson5910 (jim_thompson5910):

it's 10/3 because you go from 3y = 10 to y = 10/3 when you divide both sides by 3

OpenStudy (anonymous):

Oh, alright

OpenStudy (anonymous):

You're really good at this I may add

OpenStudy (anonymous):

Hahahaha, that was a math pun "add"

jimthompson5910 (jim_thompson5910):

lol yes a very good one

jimthompson5910 (jim_thompson5910):

thanks for the compliment

OpenStudy (anonymous):

I really do appreciate your help, you've been the best on here

jimthompson5910 (jim_thompson5910):

yw, I'm glad to be of help

OpenStudy (anonymous):

I'm good at English, Psych, and Philosophy, but I don't have much to contribute when it comes to math

jimthompson5910 (jim_thompson5910):

you'll get better at it, just keep practicing

OpenStudy (anonymous):

I know everyone has their strong suites and their weaknesses, but being enrolled in math all over again makes me feel insecure about my skills, or lack of I should say.

jimthompson5910 (jim_thompson5910):

start with what you know and build from there

jimthompson5910 (jim_thompson5910):

you'll improve your skills if you can translate it to your own language

OpenStudy (anonymous):

You have to have somewhere to start with in order to build, like the order of operations, hahaha all of these puns!

OpenStudy (anonymous):

No, but you're right. I'm only joking, but i know where I can start and build up

OpenStudy (anonymous):

My language will never be able to translate math

jimthompson5910 (jim_thompson5910):

lol fun puns (love me some puns) it will, just keep at it and you'll be able to translate

jimthompson5910 (jim_thompson5910):

try to connect it to your strengths (like eng or psych)

OpenStudy (anonymous):

Thank you for all of your help

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

There's a 100% chance i'll need more of your knowledge

OpenStudy (anonymous):

the thing is, once i feel i know how to go on with other problems< i come across on that completely doubts my ability to achieve

jimthompson5910 (jim_thompson5910):

lol alright

OpenStudy (anonymous):

ooops no backwards shark thing

jimthompson5910 (jim_thompson5910):

hmm? anyways, yes there will be problems that test your confidence, but with enough practice you'll improve your confidence

OpenStudy (anonymous):

i have to go now, but i'll be back shortly

jimthompson5910 (jim_thompson5910):

alright, cya later